# FoundationGrasp: Generalizable Task-Oriented Grasping with Foundation Models

### v1

@@ VI-B Results of Real-Robot Experiments
Task-Oriented Manipulation Experiments  This evaluation aims to prove that task-oriented grasping facilitates downstream manipulation tasks. Therefore, baselines are not evaluated here. As reported in Table V, FoundationGrasp achieves a high grasping success rate of approximately 80% on nine manipulation tasks. Specifically, in the Bottle Opening task, since the bottle cap is flat and thin, the robot attempts to grasp at the cap-neck intersection in some cases, resulting in 5 failures out of 10 trials. In terms of manipulation, despite simplifying the manipulation motions into motion primitives guided by rule-based heuristics, FoundationGrasp performs reasonably well, with an average success rate of 66.67%. Two examples of task-oriented manipulation experiments are presented in Figure 7. The performance gap between the grasping and manipulation stage is primarily due to two factors: (1) limited ability to model the relative pose between the tool and the receptacle objects; (2) absence of force feedback in tasks requiring precise contact control, such as Mug Cleaning, Ketchup Squeezing. Future work will consider extending FoundationGrasp to jointly optimize grasp poses and manipulation trajectories and fusing force modality.
@@
TABLE V: Quantitative results of task-oriented manipulation experiments
@@ VI-C Ablation Study
To gain further insights into the design choice of FoundationGrasp, we conduct two sets of ablation studies to answer the following questions:
  • How the performance of FoundationGrasp is affected by semantic and geometric branches?
@@
  • How do object and task knowledge help the generalization to novel elements out of the training set?
@@
Ablation on Semantic and Geometric Branches  In this study, we compare FoundationGrasp to three ablated versions: (1) with semantic branch only (i.e., w/ Sem); (2) with geometric branch only (i.e., w/ Geo); (3) vanilla model (i.e., w/o Sem & Geo). The result is presented in Figure 9 (a). Here, we only show mAPs of corresponding held-out settings. For all held-out settings, the full model achieves the best performance, while the vanilla model performs the worst. This demonstrates the effectiveness of semantic-geometric branch architecture. Specifically, for the held-out instance setting, where objects only have limited variances in terms of geometric and visual features, geometric knowledge is essential for unifying such variances. The relations between tasks are easier to capture when tasks are abstracted as linguistic concepts. Therefore, semantic knowledge contributes more to relating different tasks in the held-out task setting. Semantic and geometric branches play an equally important role when generalizing to novel object classes.
@@
Ablation on Object and Task Knowledge  In this study, we compare FoundationGrasp to three ablated versions: (1) with object knowledge only (i.e., w/ Obj); (2) with task knowledge only (i.e., w/ Task); (3) vanilla model (w/o Obj & Task). The result is presented in Figure 9 (b). It is clear from held-out class and task settings that object and task knowledge are essential for generalizing to novel object classes and tasks, respectively. This aligns with the intuition. The full model only outperforms w/ Obj by 1.89% in the held-out class setting. For the held-out instance setting, w/ Obj and w/ Task achieve similar performances (88.47% and 86.98%). Overall, incorporating both object and task knowledge outperforms all ablated versions.
@@
Figure 9: Ablation study: (a) ablation on semantic and geometric branches; (b) ablation on object and task knowledge.
@@ VI-D User Study
To deploy FoundationGrasp on household robots operating in everyday life, it is also necessary to study the practical applicability from the user perspective (e.g., whether the behavior of FoundationGrasp aligns with user needs, preferences, safety, and comfort). Therefore, we conduct a user study with 10 participants on 10 test objects selected from the real-robot experiments benchmark. Once all objects are tested, participants are invited to fill out a questionnaire with Likert scale. Participants are also encouraged to share their experiences or describe any problems they encounter during the user study.
Table VI reports the objective metrics of the user study. On average, the robot can successfully grasp each object with less than two attempts. Figure 10 presents the subjective responses of participants to the questionnaire. Average ratings in all dimensions are above 3 (i.e., Agree or Strongly Agree). Most participants agree that the robot is aware of the task-oriented effect and is able to grasp objects in a task-oriented manner. Some comment that FoundationGrasp exhibits high repeatability (i.e., generates grasp poses around the same point) given the same object and task with different expressions. Participants also note that FoundationGrasp is robust to arbitrary object poses. Such a desirable property is due to the data augmentation performed during training. Regarding safety, one participant expresses concerns about potential collisions with users, resulting in a disagreement on the third statement. Overall, participants suggest the system could be improved in the following ways: (1) The system should consider the state of the object. For example, the generated grasps should be far away from the body of the pot if the pot is hot. (2) The system should be able to understand language instructions directly specifying the part (e.g., “grasp the handle of the pot”).
@@
TABLE VI: Objective evaluation of user study
VII Discussions and Limitations
@@
Despite compelling results, we observe typical failure modes or limitations during experiments. This section discusses these limitations from data, algorithm, and hardware perspectives and provides insights into future work.
@@
Grasp Data Acquisition  While FoundationGrasp achieves superior generalization performance compared to previous works, the data collection phase still requires labor-intensive manual annotation. Besides, some annotations do not align well with how humans perform task-oriented grasping, which may further hurt the performance. This point can be reflected by the neutral attitude of several participants towards the statement “The robot was able to perform grasping in human-like behavior” in the user study. One potential solution is to learn from human demonstrations. Previous works have proved that human demonstrations (e.g., images or videos) encode constraints and semantics of human behaviors for robot skill learning. Learning with general-purpose simulators of the physical world is also an exciting direction. Recent advances in AIGC (e.g., Sora[46]) open up the potential for generating synthetic demonstration data without a human in the loop.
@@
Sampling-and-Evaluation Strategy  FoundationGrasp follows a two-stage, sampling-and-evaluation pipeline. Using the off-the-shelf task-agnostic grasp sampler reduces the complexity of constructing a TOG pipeline. However, such a design also leads to two problems. First, we assume all grasp candidates from the sampling phase are stable and treat them equally in the evaluation phase. However, we observe that FoundationGrasp tends to bias marginal grasp poses, for example, near the end of the handle. Although functionally correct, these grasp poses are not configurationally robust to perturbations, such as unexpected contact or calibration error. This would lead to potential grasping failure. Second, the first stage samples grasps uniformly over the point cloud. However, robots would only interact with specific regions of an object. Exhaustively evaluating every candidate grasp significantly increases the computational cost of FoundationGrasp and hinders real-time prediction. Future work plans to merge task-agnostic sampling with task-oriented evaluation into a unified framework.
@@
Robot Gripper Generalization  Regarding hardware setup, we currently only consider the parallel jaw gripper. Therefore, FoundationGrasp cannot transfer to other types of grippers, such as three-finger grippers and five-finger dexterous hands. Although such a methodology can be generally applied to training any gripper-dependent TOG model, it is still far from a universal TOG method in terms of both software and hardware. Ideally, the algorithm would take both object geometry and gripper attributes as inputs and output the grasp poses or contact points. The long-term goal is to develop a universal TOG method that can be readily applied to any household robot with any type of gripper.
@@ VIII Conclusion
In this paper, we present FoundationGrasp, a foundation model-based TOG framework. Compared to previous works, which limit the prior knowledge to a closed-set scope, FoundationGrasp leverages the open-ended knowledge from foundation models to learn generalizable TOG skills. Evaluation on the LaViA-TaskGrasp dataset demonstrates the superiority of FoundationGrasp over existing TOG methods when generalizing to novel object instances, object classes, and tasks. Furthermore, the effectiveness of FoundationGrasp is validated in task-oriented grasping and manipulation experiments on a 7 DoF robotic arm. Lastly, limitations and potential solutions are thoroughly discussed to provide venues for future work.

### v2

@@ VI-B Results of Real-Robot Experiments
Task-Oriented Manipulation Experiments  This evaluation aims to prove that task-oriented grasping facilitates downstream manipulation tasks. Therefore, baselines are not evaluated here. As reported in Table VI, FoundationGrasp achieves a high grasping success rate of approximately 80% on nine manipulation tasks. Specifically, in the Bottle Opening task, since the bottle cap is flat and thin, the robot attempts to grasp at the cap-neck intersection in some cases, resulting in 5 failures out of 10 trials. In terms of manipulation, despite simplifying the manipulation motions into motion primitives guided by rule-based heuristics, FoundationGrasp performs reasonably well, with an average success rate of 66.67%. Two examples of task-oriented manipulation experiments are presented in Figure 10. The performance gap between the grasping and manipulation stages is primarily due to two factors: (1) limited ability to model the relative pose between the tool and the receptacle objects; (2) absence of force feedback in tasks requiring precise contact control, such as Mug Cleaning and Ketchup Squeezing. Future work will consider extending FoundationGrasp to jointly optimize grasp poses and manipulation trajectories and incorporate force feedback.
In summary, there are three primary challenges encountered in real-robot experiments:
  • Limited Generalization: A significant challenge is the limited ability to generalize to novel objects and tasks unseen during training. FoundationGrasp addresses this by leveraging open-ended knowledge from foundation models, thereby enhancing the generalization capabilities beyond those of previous TOG methods.
  • Occlusion: Another challenge arises from occlusions. If the target object or functional part is significantly occluded, the robot may struggle to generate valid grasp candidates or may produce false evaluations.
  • Planning: The third challenge involves planning. Since FoundationGrasp only focuses on generating TOG poses for the end-effector, the off-the-shelf planner is utilized to move the gripper to the target pose, during which planning failures might occur.
@@
TABLE VI: Quantitative results of task-oriented manipulation experiments
@@ VI-C Ablation Study
To gain further insights into the design choices of FoundationGrasp, we conduct a series of ablation studies to answer the following questions:
  • How is the performance of FoundationGrasp affected by the semantic and geometric branches?
@@
  • How do object and task knowledge contribute to the generalization to novel elements outside the training set?
  • How do different fusion strategies in the task-oriented grasp evaluator affect the performance of FoundationGrasp?
  • How does the choice of a VLM in the geometric knowledge encoder impact the performance of FoundationGrasp?
  • How does the choice of an LLM in the semantic knowledge encoder impact the performance of FoundationGrasp?
  • How does the choice of an LLM in language description generation impact the performance of FoundationGrasp?
@@
Ablation on Semantic and Geometric Branches  In this study, we compare FoundationGrasp to three ablated versions: (1) with the semantic branch only (i.e., w/ Sem); (2) with the geometric branch only (i.e., w/ Geo); (3) the vanilla model (i.e., w/o Sem & Geo). The results are presented in Figure 11 (a). Here, we only show mAPs of the corresponding held-out settings. For all held-out settings, the full model achieves the best performance, while the vanilla model performs the worst. This demonstrates the effectiveness of the semantic-geometric branch architecture. Specifically, for the held-out instance setting, where objects have limited variations in terms of geometric and visual features, geometric knowledge is essential for unifying these variances. The relations between tasks are easier to capture when tasks are abstracted as linguistic concepts, making semantic knowledge more critical in relating different tasks in the held-out task setting. To conclude, both semantic and geometric branches play equally important roles when generalizing to novel object classes.
@@
Ablation on Object and Task Knowledge  In this study, we compare FoundationGrasp to three ablated versions: (1) with object knowledge only (i.e., w/ Obj); (2) with task knowledge only (i.e., w/ Task); (3) the vanilla model (i.e., w/o Obj & Task). The results are presented in Figure 11 (b). It is clear from the held-out class and task settings that object and task knowledge are crucial for generalizing to novel object classes and tasks, respectively. This finding aligns with our intuition. In the held-out class setting, the full model only outperforms w/ Obj by 1.89%. For the held-out instance setting, w/ Obj and w/ Task achieve similar performances (88.47% and 86.98%). Overall, incorporating both object and task knowledge leads to superior performance compared to all ablated versions.
@@
Figure 11: (a) Ablation study on semantic and geometric branches. (b) Ablation study on object and task knowledge.
Since each of the following ablation studies shares similar conclusions across the three held-out settings, we only report the results of the held-out class setting for simplicity.
Ablation on Fusion Strategy   In this study, we investigate how different fusion strategies impact the performance of FoundationGrasp. Specifically, we compare the original cross-attention-based fusion (denoted as Cross-Attn) with the following alternative methods: (1) double attention [56] (denoted as Double-Attn), which replaces the cross attention in each transformer block with two consecutive attention blocks for feature gathering and distribution; (2) spatial-channel attention [57] (denoted as SC-Attn), which uses spatial and channel attention in parallel; (3) feature-wise linear modulation [58] (denoted as FiLM), which modifies features via a learnable affine transformation based on conditioning information; and (4) concatenation (denoted as Concat), which concatenates all input features channel-wise in each transformer block without any learnable parameters. As shown in Table VII, the simple concatenation fusion achieves the worst performance, falling behind Cross-Attn by 13.51%, 13.75%, and 10.42% across the three evaluation metrics. Among the attention-based fusion strategies, SC-Attn demonstrates a slight advantage over Cross-Attn, with improvements of 0.75% in Instance mAP and 0.85% in Class mAP. Despite these marginal gains, we use Cross-Attn due to its balance between performance and computational cost.
Method Held-out Class Performance (%)
Instance mAP Class mAP Task mAP
Cross-Attn 88.74 86.09 84.33
Double-Attn 85.43 83.43 80.78
SC-Attn 89.49 86.94 84.01
FiLM 83.87 80.54 81.29
Concat 75.23 72.34 73.91
TABLE VII: Ablation study on fusion strategy
Ablation on LLMs in Semantic Knowledge Encoder  In this study, we equip FoundationGrasp with four different LLMs, including BERT, RoBERTa[59], LLaMA 3[60], and T5[61], in the semantic knowledge encoder and compare their performances. The results are detailed in Table VIII. Overall, all LLMs perform comparably, with LLaMA 3 showing a slight edge with mAPs of 89.63%, 87.53%, and 86.48%. The consistent performance across the LLMs is due to their extensive pre-training on large text corpora, which endows them with a robust understanding of commonsense semantic knowledge. Despite the similarities, we select Google pre-trained BERT as the semantic knowledge encoder in FoundationGrasp, balancing effective knowledge understanding with computational efficiency.
Method Held-out Class Performance (%)
Instance mAP Class mAP Task mAP
BERT 88.74 86.09 84.33
RoBERTa 87.54 86.45 84.09
LLaMA 3 89.63 87.53 86.48
T5 88.73 87.90 85.34
TABLE VIII: Ablation study on LLMs in the semantic knowledge encoder
Ablation on VLMs in Geometric Knowledge Encoder  In this study, we explore the impact of different VLMs by substituting the CLIP model used in the geometric knowledge encoder with three alternatives: (1) BLIP-2 [62], noted for its strong performance in multi-modal tasks like Visual Question Answering; (2) LLaVa [63], which combines a CLIP vision encoder with a LLaMA language encoder; and (3) CLIP-V, which removes the CLIP language encoder. The results, summarized in Table IX, show that all three VLMs perform comparably, with CLIP achieving the highest performance in Class mAP (86.09%) and Task mAP (84.33%). Notably, CLIP-V, which relies solely on visual inputs, underperforms relative to CLIP by 5.21%, 4.26%, and 3.35% across the evaluation metrics. This finding emphasizes the significance of multi-modal knowledge for effective geometric understanding, as supported by previous studies [64, 65].
Method Held-out Class Performance (%)
Instance mAP Class mAP Task mAP
CLIP 88.74 86.09 84.33
BLIP-2 86.49 84.92 82.09
LLaVa 88.98 85.43 84.01
CLIP-V 83.53 81.83 80.98
TABLE IX: Ablation study on VLMs in the geometric knowledge encoder
Ablation on LLMs in Language Description Generation  In this study, we evaluate the performance of FoundationGrasp using two different LLMs for language description generation: GPT-4 and GPT-3.5. As expected, the more advanced GPT-4 slightly outperforms GPT-3.5 across the three evaluation metrics by 0.76%, 1.7%, and 1.24%, respectively. However, the performance differences between the two models are relatively minor.
Method Held-out Class Performance (%)
Instance mAP Class mAP Task mAP
GPT-4 88.74 86.09 84.33
GPT-3.5 87.98 84.39 83.09
TABLE X: Ablation study on LLMs for language description generation
To summarize, the key insights from the ablation studies on foundation models are two-fold:
  • FoundationGrasp demonstrates robust performance across various VLMs and LLMs, with only minor performance differences between different foundation models.
  • The modular design of FoundationGrasp allows for the seamless integration of any state-of-the-art foundation models. This adaptability ensures that as foundation models evolve, the performance of FoundationGrasp will also improve accordingly.
@@ VI-D User Study
To deploy FoundationGrasp on household robots operating in everyday life, it is also necessary to study its practical applicability from the user perspective (e.g., whether the behavior of FoundationGrasp aligns with user needs, preferences, safety, and comfort). Therefore, we conduct a user study with 10 participants on 10 test objects selected from the real-robot experiments benchmark. Once all objects are tested, participants are invited to fill out a questionnaire with a Likert scale. Participants are also encouraged to share their experiences or describe any problems they encounter during the user study.
Table XI reports the objective metrics of the user study. On average, the robot can successfully grasp each object with less than two attempts. Figure 12 presents the subjective responses of participants to the questionnaire. Average ratings in all dimensions are above 3 (i.e., Agree or Strongly Agree). Key findings from the user study include: (1) Most participants agree that the robot is aware of the task-oriented effect and can grasp objects in a task-oriented manner; (2) Some comment that FoundationGrasp exhibits high repeatability, generating grasp poses around the same point for the same object and task with different expressions; (3) Participants find that FoundationGrasp is robust to arbitrary object poses, attributing this to the data augmentation performed during training. Regarding safety, one participant expresses concerns about potential collisions with users, resulting in a disagreement on the third statement. Overall, participants suggest the system could be improved in the following ways: (1) The system should consider the state of the object. For example, the generated grasps should be far away from the body of the pot if the pot is hot. (2) The system should be able to understand language instructions directly specifying the part (e.g., “grasp the handle of the pot”).
@@
TABLE XI: Objective evaluation of user study
VII Limitations and Future Work
@@
The results of perception experiments verify the claim that FoundationGrasp outperforms existing TOG methods when generalizing to novel object instances, object classes, and tasks out of the training set. Meanwhile, real-robot experiments have also demonstrated the effectiveness of FoundationGrasp in real-world applications. Despite compelling results, we observe typical failure modes or limitations during experiments. This section discusses these limitations from data, algorithm, and hardware perspectives and provides insights into future work.
@@
Grasp Data Acquisition  While FoundationGrasp achieves superior generalization performance compared to previous works, the data collection phase still requires labor-intensive manual annotation. Additionally, some annotations do not align well with how humans perform task-oriented grasping, which may further hurt performance. This issue is reflected by the neutral attitude of several participants towards the statement “The robot was able to perform grasping in human-like behavior” in the user study. One potential solution is to learn from human demonstrations. Previous works have shown that human demonstrations (e.g., images or videos) encode constraints and semantics of human behaviors for robot skill learning. Learning with general-purpose simulators of the physical world is also an exciting direction. Recent advances in AIGC (e.g., Sora [66]) open up the potential for generating synthetic demonstration data without a human in the loop.
@@
Sampling-and-Evaluation Pipeline  FoundationGrasp currently follows a two-stage, sampling-and-evaluation pipeline. Using the off-the-shelf task-agnostic grasp sampler reduces the complexity of constructing a TOG pipeline. However, such a design also leads to two problems. First, we assume all grasp candidates from the sampling phase are stable and treat them equally in the evaluation phase. However, we observe that FoundationGrasp tends to bias marginal grasp poses, such as grasps near the end of the handle. Although functionally correct, these grasp poses are not configurationally robust to perturbations, such as unexpected contact or calibration error. This would lead to potential grasping failure. Second, the first stage samples grasps uniformly over the point cloud. However, robots would only interact with specific regions of an object. Exhaustively evaluating every candidate grasp significantly increases the computational cost of FoundationGrasp and hinders real-time prediction. Future work plans to merge task-agnostic sampling with task-oriented evaluation into a unified framework.
@@
Robot Gripper Generalization  Regarding hardware setup, we currently only consider the parallel jaw gripper. Therefore, FoundationGrasp cannot be directly transferred to other types of grippers, such as three-finger grippers and five-finger dexterous hands. Although such a methodology can be generally applied to training any gripper-dependent TOG model, it is still far from a universal TOG method in terms of both software and hardware. Ideally, the algorithm would take both object geometry and gripper attributes as inputs and output the grasp poses or contact points. The long-term goal is to develop a universal TOG method that can be readily applied to any household robot with any type of gripper.
@@ VIII Conclusion
In this paper, we present FoundationGrasp, a foundation model-based TOG framework. Compared to previous works, which limit the prior knowledge to a closed-set scope, FoundationGrasp leverages the open-ended knowledge from foundation models to learn generalizable TOG skills. Evaluation on the LaViA-TaskGrasp dataset demonstrates the superiority of FoundationGrasp over existing TOG methods when generalizing to novel object instances, object classes, and tasks. Furthermore, the effectiveness of FoundationGrasp is validated in task-oriented grasping and manipulation experiments on a 7-DoF robotic arm. Lastly, limitations and potential solutions are thoroughly discussed to provide venues for future work.